OpenStudy (anonymous):

What is the quotient of 7^-1/7^-2? A. 1/343 B. 1/7 C. 7 D. 49

OpenStudy (anonymous):

Please someone help me

Nnesha (nnesha):

hint: \[\huge\rm x^{-2} = \frac{ 1 }{ x^2 }\]\ if there is negative exponent then u should flip the fraction

OpenStudy (anonymous):

Ok

Nnesha (nnesha):

and when there are same bases at the denominator and at the numerator u can move exponent from bottom to the top for example \[\huge\rm \frac{ x^m }{ x^n} = x^{m-n}\]